OpenStudy (anonymous):

There are two consecutive even numbers such that six times the smaller number minus five times the greater number is -2. What is the smaller number?

OpenStudy (anonymous):

X = smaller number, x+2 = larger number 6x-5(x+2)=-2 6x-5x-10=-2 x=8